A stone is found to weigh 6.2 in air, 4.5 in water, and 5.3 N in an unknown liquid. Find a.) the density of the stone and the unknown liquid, b.) the buoyant force of the liquid on the stone, c.) the buoyant force of water on the stone, and d) the volume of the stone.


1
Expert's answer
2021-10-25T15:23:54-0400

"P_{air} = 6.2N"

"P_{water}= 4.5N"

"P_{unliquid}= 5.3N"

"g= 9.8\frac{m}{s^2}"

"\text{in water:}"

"\rho_{water} = 1000\frac{kg}{m^3}"

"BF_{water} = P_{air}-P{water}= 6.2-4.5= 1.7N"

"BF_{water} = \rho_{water}gV_{stone}"

"V_{stone} = \frac{BF_{water}}{ \rho_{water}g}=\frac{1.7}{1000*9.8}\approx 1.73*10^{-4}m^3"

"\text{in air : }"

"P_{air} = m_{stone}g"

"\rho_{stone}= \frac{m_{stone}}{V_{stone}}=\frac{P_{air}}{gV_{stone}}"

"\rho_{stone}= \frac{6.2}{9.8*1.73*10^{-4}}=3.66*10^{3}\frac{kg }{m^3}"

"\text{in an unknown liquid:}"

"BF_{unliquid}= P_{air} -P_{unliquid}= 6.2-5.3 = 0.9N"

"BF_{unliquid} = \rho_{unliquid}gV_{stone}"

"\rho_{unliquid} =\frac{ BF_{unliquid}} {gV_{stone}}=\frac{0.9}{9.8*1.73*10^{-4}}=0.53*10^3\frac{kg}{m^3}"


"\text{Answer:}"

"a)\rho_{stone}=3.66*10^{3}\frac{kg }{m^3}"

"\rho_{unliquid} =0.53*10^3\frac{kg}{m^3}"

"b)BF_{unliquid}= 0.9N"

"c)BF_{water} = 1.7N"

"d)V_{stone} = 1.73*10^{-4}m^3"